If you pay $4 for a call option on JPM stock with an at the money strike price of $100 and at the same time you write a call option with a strike price of $110 for which you receive $1; how much money do you make or lose on the transaction if the stock goes to either $120; $105; or $90?
Answer #1

Buy $100 strike call option for $4 and write $110 strike call option for $1. This is a debit call spread.

X1 = 100

X2 = 110

Net premium paid = 4 - 1 = $3

Profit = max(St - X1, 0) - max(St - X2, 0) - Net premium paid.

1) St = 120

Profit = max(120 - 100, 0) - max(120 - 110, 0) - 3

Profit = 20 - 10 - 3

Profit = $7

2) St = 105

Profit = max(105 - 100, 0) - max(105 - 110, 0) - 3

Profit = 5 - 0 - 3

Profit = $2

3) St = 90

Profit = max(90 - 100, 0) - max(90 - 110, 0) - 3

Profit = 0 - 0 - 3

Profit = -$3 Or a loss of $3